OpenStudy (anonymous):

Use the change of base formula to evaluate log3(73). I am unsure of how to solve this... 0.0 (Metal and fan)

OpenStudy (tylerd):

3^x = 73 xlog(3)=log(73) x=log(73)/log(3)

OpenStudy (tylerd):

dont know change of base formula but thats how id solve

OpenStudy (anonymous):

wait i dont understand..are you dividing 73 by 3?

OpenStudy (tylerd):

im taking log of both sides

OpenStudy (tylerd):

and when you take log of something to the power x, the x drops down is is multiplied

OpenStudy (tylerd):

im dividing log(73) by log(3)

OpenStudy (anonymous):

Hmmm ok well my answer choices are 2.641 4.290 1.863 3.905

OpenStudy (anonymous):

and that does not match up

OpenStudy (tylerd):

it does

OpenStudy (anonymous):

oh ok let me try again

OpenStudy (tylerd):

\[\frac{ \log(73) }{ \log(3) }=3.905\]

OpenStudy (anonymous):

Ok, i see that. Hmmm...for some reason when i divide i get 24.333